Stock XYZ has a current dividend of $2.00
The dividend will grow for 3 periods at a rate of 4%
From period 3 onward , dividends will grow at a rate of 1%
Based on the riskiness for XYZ company, the discount rate for its stock is 6.00%

Given the information provided, what is the current price of XYZ stock ?

A

$46.13

B

$38.16

C

$45.44

D

$43.93

please show work
Answer #1

D1=(2*1.04)=2.08

D2=(2.08*1.04)=2.1632

D3=(2.1632*1.04)=2.249728

Value after year 3=(D3*Growth rate)/(Discount rate-Growth rate)

=(2.249728*1.01)/(0.06-0.01)

=45.4445056

Hence current price=Future dividend and value*Present value of discounting factor(rate%,time period)

=2.08/1.06+2.1632/1.06^2+2.249728/1.06^3+45.4445056/1.06^3

=$43.93(Approx).